A man has a daughter and a son.
The son is three years older than the daughter.
In 17 years the son will be twice as old as the daughter is now.
What is the daughter's present age?

Answer by Paul(988)

Son's age = 3+x
Daugther's age = x
In 17 years son's age = x+17+3-->x+20

Equation:
2(x)=x+20
2x=x+20
x=20

20+3=23
Hence, the age for the son is 23 years and the age for the dauther is 20 years.
Paul.